OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -9 - 5t. Find the instantaneous velocity at t = 4 by finding the derivative.

OpenStudy (anonymous):

please help!!!!

OpenStudy (aum):

s(t) = -9 - 5t can you find ds/dt ?

OpenStudy (anonymous):

no im lost on this particular problem..im stuck :-(

OpenStudy (aum):

You have no calculus background? You don't know the derivative of say, x^3, with respect to x?

OpenStudy (anonymous):

yes i do. but at this point, i'm basically going back and reviewing this stuff, i haven't looked at it in a while.

OpenStudy (aum):

What is the derivative of a constant?

OpenStudy (anonymous):

wait i think i have the answer

OpenStudy (anonymous):

is it -5

OpenStudy (anonymous):

-5t

OpenStudy (anonymous):

s'(t)= -5

OpenStudy (aum):

\[ s(t) = -9 - 5t \\ s'(t) = -5 \\ s'(4) = -5 \]

OpenStudy (aum):

The instantaneous velocity is s'(t) which is -5 and it is independent of time. So the velocity remains the same at -5 no matter what the value of t. Here they want instantaneous velocity at t = 4. It is the same value of -5.

OpenStudy (anonymous):

thanks!!!have time for one more?

OpenStudy (anonymous):

Find an explicit rule for the nth term of the sequence. 3, -12, 48, -192, ... @aum

OpenStudy (aum):

Do you see a common pattern in the numbers? What can you do to the first number to get the second number? Can the same thing be done to the second number to get the third number? Can the same thing be done to the third number to get the fourth number?